OpenStudy (anonymous):

answer in radical form 16^(1/3)y^3

OpenStudy (anonymous):



OpenStudy (saifoo.khan):

\[y^3\sqrt[3]{16}\]

OpenStudy (anonymous):

the y^3 is not under a radical because the denominator of its exponent is a 1; the exponent of 1/3 means cube root; the factors to 8*2 where 8 is a perfect cube \[16^{1/3}y^3=y^3\sqrt[3]{8*2}=2y^3\sqrt[3]{2}\]

OpenStudy (anonymous):

the original question was